OpenStudy (anonymous):

Find the Limit. Use l'hospital's rule where appropriate. lim as x goes to 0 from the right of (ln x / x) I did l'hospital and got... lim as x goes to 0 from the right of (1/x)/1, but that is still indeterminate so i did it again, but the denominator would = 0 so we can't divide by 0 the book answer is - infinity, but i dont understand why

OpenStudy (anonymous):

0 = cos(x) -sin(x) cos(x) = sin(x) x = 45, 225

OpenStudy (tkhunny):

One thing you might want to keep in mind, even though it will not speed this problem on its way. \[\sin(x) + \cos(x) = \sqrt{2}\cdot\cos\left(x-\frac{\pi}{4}\right)\] Just keep it somewhere in the back of your head. You may need the general form, some day.

OpenStudy (anonymous):

All right, so: \[ \lim_{x\to0^+}\left(\frac{\ln(x)}{x}\right)=\lim_{x\to0^+}\left(\frac{\frac{d}{dx}\ln(x)}{\frac{d}{dx}x}\right)=\lim_{x\to0^+}\left(\frac{\frac{1}{x}}{1}\right)=\lim_{x\to0^+}\left(\frac{1}{x}\right)=\infty \]Are you sure it's not from the left?

OpenStudy (anonymous):

Yes, I'm sure. So... in your answer... since x is not ACTUALLY 0 (just very close) we get infinity because it's 1/(a very very small number)?